A boat takes 8 hours to travel 80 miles upstream and 5 hours to return to its starting point. Find the speed of the boat in still water and the speed of the current.
Answer by checkley71(8403)

80/8=10MPH UPSTREAM & 80/5=16MPH THE SPEED OF THE BOAT IS (16+10) /2 OR 13MPH &
THE CURRENT IS 3MPH
PROOF 13-3=10=80/8=10 & 13+3=16=80/5=16